How many formula units of salt are contained in 1 teaspoon of salt? How many chlorine atoms would the sample contain? 1 teaspoon NaCl = 8.358g
1
Expert's answer
2018-09-27T05:48:56-0400
If by formula units you mean moles, then you simply take the mass of the NaCl and this mass is regarded as the molar mass (g/mol). Next, use this mass (58.42g/mol) and use stoichiometry.

n(NaCl) = m/M = 8.358g/58.42g/mol = 0.143 mol

N(NaCl) = n*NA = 0.143mol*6.02·1023mol-1 = 8.6·1022 molecules

N(Cl) = 8.6·1022 atoms